A CYLINDER HAS A RADIUS OF 5IN. IF THE VOLUME OF THE CYLINDER IS 250IN^3, WHAT IS THE HEIGHT OF THE CYLINDER? (HINT THE VOLUME OF A CYLINDER IS GIVEN BY V=R^2H)..THANKS IN ADVANCE
Answer by stanbon(75887)

A CYLINDER HAS A RADIUS OF 5IN. IF THE VOLUME OF THE CYLINDER IS 250IN^3, WHAT IS THE HEIGHT OF THE CYLINDER? (HINT THE VOLUME OF A CYLINDER IS GIVEN BY V=R^2H)..
-------------
V = R^2 H
250 = 5^2 H
250 = 25 H
H=10 in
Cheers,
Stan H.